OpenStudy (anonymous):

I am needing help! Find the volume of a box that measures 4 1/3 feet by 1 1/2 feet by 3 1/3 feet

OpenStudy (jdoe0001):

hmmm ok.... so... \(\bf \bf 4\frac{1}{2}\times 1\frac{1}{2}\times 3\frac{1}{3}\) you would want to make the "mixed fraction" to a so-called proper one first...so you're left with only one numerator and one denominator

OpenStudy (jdoe0001):

\(\bf 4\frac{1}{2}\times 1\frac{1}{2}\times 3\frac{1}{3}\implies \cfrac{4\cdot 2+1}{2}\times \cfrac{1\cdot 2+1}{2}\times \cfrac{3\cdot 3+1}{3}\)

OpenStudy (anonymous):

13 x 3 x 10 = 390? @jdoe0001

OpenStudy (anonymous):

390 over 6?

OpenStudy (jdoe0001):

hmmm well.. not quite... what would you get for hte fractions anyway?

OpenStudy (anonymous):

9/2 x 3/2 x 10/ 3

OpenStudy (anonymous):

270?

OpenStudy (anonymous):

the equation is 4 1/3 not 4 1/2 so then it would be 13- correct

OpenStudy (jdoe0001):

yeap \(\bf 4\frac{1}{2}\times 1\frac{1}{2}\times 3\frac{1}{3}\implies \cfrac{4\cdot 2+1}{2}\times \cfrac{1\cdot 2+1}{2}\times \cfrac{3\cdot 3+1}{3} \\ \quad \\ \cfrac{9}{2}\times \cfrac{3}{2}\times \cfrac{10}{3}\implies \cfrac{9\times 3\times 10}{2\times 2\times 3}\implies \cfrac{270}{12} \\ \quad \\ \textit{then we divide top and bottom by }6 \\ \quad \\ \cfrac{\cancel{ 270 }}{\cancel{ 12 }}\implies \cfrac{45}{2}\implies 22\frac{1}{2}\)

OpenStudy (jdoe0001):

ohh ahemm... just read you... ok so

OpenStudy (anonymous):

thank you!!!!

OpenStudy (jdoe0001):

\(\bf 4\frac{1}{3}\times 1\frac{1}{2}\times 3\frac{1}{3}\implies \cfrac{4\cdot 3+1}{3}\times \cfrac{1\cdot 2+1}{2}\times \cfrac{3\cdot 3+1}{3} \\ \quad \\ \cfrac{13}{3}\times \cfrac{3}{2}\times \cfrac{10}{3}\implies \cfrac{13\times 3\times 10}{3\times 2\times 3}\implies \cfrac{390}{18} \\ \quad \\ \textit{then we divide top and bottom by }6 \\ \quad \\ \cfrac{\cancel{ 390 }}{\cancel{ 18 }}\implies \cfrac{65}{3}\implies 21\frac{2}{3}\)